### v1

# FT2TF: First-Person Statement Text-To-Talking Face Generation

Abstract
Talking face generation has gained immense popularity in the computer vision community, with various applications including AR/VR, teleconferencing, digital assistants, and avatars. Traditional methods are mainly audio-driven ones which have to deal with the inevitable resource-intensive nature of audio storage and processing. To address such a challenge, we propose FT2TF – First-Person Statement Text-To-Talking Face Generation, a novel one-stage end-to-end pipeline for talking face generation driven by first-person statement text. Moreover, FT2TF implements accurate manipulation of the facial expressions by altering the corresponding input text. Different from previous work, our model only leverages visual and textual information without any other sources (e.g. audio/landmark/pose) during inference. Extensive experiments are conducted on LRS2 and LRS3 datasets, and results on multi-dimensional evaluation metrics are reported. Both quantitative and qualitative results showcase that FT2TF outperforms existing relevant methods and reaches the state-of-the-art. This achievement highlights our model’s capability to bridge first-person statements and dynamic face generation, providing insightful guidance for future work.
Under Review
1 Introduction
Talking face generation, with the goal of creating a sequence of faces to simulate the appearance of a person talking, has surged in significance amid the metaverse boom and technological advancements. This pursuit has gained substantial popularity and importance within the computer vision community, influencing applications including augmented reality [16], virtual reality [9], computer games [30, 29], teleconferencing [7, 3], video dubbing [11, 18], movie special effects [43], and the development of digital assistants [33] and avatars [9, 49].
Traditional methods for talking face generation mainly rely on audio input, which leverages visual and audio cues. These methods, exemplified by works like [28, 49, 48, 17, 46], show impressive performance. However, they come with a significant drawback: the need for high-quality audio, leading to strict requirements on storage space, data transmission, and recording environments, as highlighted in [34]. Addressing such challenges holds significant value not only for the scientific community but also for the industry, aiming to mitigate the environmental impact of AI technologies  [32]. Consequently, one crucial question has arisen: Is it feasible to substitute audio with text inputs while ensuring detailed facial expressions of generated frames?
Figure 1: High-quality text-driven talking face generation. FT2TF aims to generate realistic talking faces using two inputs: (i) reference talking face frames and (ii) first-person statement text.
There exist two primary types of text-driven face generation methods. The first involves using third-person text descriptions where a facial expression is described from an external perspective. These descriptions are leveraged to generate realistic faces using end-to-end GAN architectures [13, 25, 41]. The second path employs first-person statements which are the speaker’s spoken captions captured in text form. While this approach aligns more closely with talking face generation, it has not been thoroughly explored. Previous methods mainly conduct a two-stage framework including text-to-speech and speech-to-face procedures [27, 45]. These approaches generate audio waveforms from text statements and further generate faces based on synthesized audio [31]. Therefore, methods that follow a one-stage end-to-end pattern to generate talking faces driven by first-person statements remain unestablished. Based on this, our objective is to revolutionize the synthesis of coherent and natural talking face videos through first-person statements, encompassing well-synchronized lip movements, accurate cheek actions, detailed textures, and overall facial expressions, as illustrated in Figure 1.
To overcome the challenges above, we propose First-Person Statement Text-To-Talking Face Generation (FT2TF) – a one-stage end-to-end pipeline that effectively generates high-quality talking faces from visual-textual cross-modality. Different from previous work, our pipeline encapsulates the entire face generation process without any other sources (e.g. audio, landmarks, poses) being used during inference.
In summary, our contribution is threefold:
  1. 1.
    End-to-end Continuous-Frame Talking Face Generation. We propose FT2TF – an effective one-stage pipeline that generates realistic talking faces by integrating visual and textual input. FT2TF evidently enhances face naturalness, synchronizes lip movements, and improves detailed facial expressions.
  2. 2.
    Effective Facial Expression Manipulation. FT2TF enables manipulations of the speaker’s facial expressions and lip movements by altering corresponding input text.
  3. 3.
    State-of-the-Art Performance. We conduct extensive experiments on the LRS2 [1] and LRS3 [2] datasets with both qualitative and quantitative results to demonstrate the effectiveness of FT2TF. This strongly proves that our model can consistently generate high-quality talking faces across different conditions.
2 Related Work
2.1 Audio-Driven Talking Face Generation
The majority of the work on talking face generation is audio-driven which needs both audio and video ground truth as the input, focusing on lip region/whole face generation.
Lip Generation. As a groundwork, Chen et al. proposes ATVGnet [6] that uses audio signal and reference image to generate the region around the lip, while the remaining parts of the face remain static from frame to frame. Based on that, Wav2Lip [28] generates dynamically changed faces through lip masking and reconstruction strategy. Following the previous lip masking strategy, IP_LAP [46] works as state-of-the-art by using a Transformer-based [36] encoder to predict the landmarks for further lip generation. Although these methods generate high-quality dynamic faces, all of them only focus on lip generation, failing to capture facial expression details of the entire face.
Whole Face Generation. Zhou et al. [49] disentangles audio sources into content and speaker embeddings, followed by an LSTM-based module for landmark prediction to generate the talking head. Based on that, models that use input sources other than audio have been proposed. PC-AVS [48] utilizes the pose source to control the head pose of the synthesized faces, while EAMM [17] uses pose sequence, emotion source, and audio ground truth to generate the target video manipulated by different emotion guidances. In contrast to existing methods that involve multiple input resources for face manipulation, our model only uses text and reference face frames as the input source. This ensures the efficient generation of high-quality facial representations, free from input requirements such as constraints of audio quality.
Figure 2: Overview of FT2TF Pipeline. The FT2TF pipeline employs two specialized Text Encoders, the Global Emotion Text Encoder, and the Linguistic Text Encoder, for extracting emotional and linguistic text features, respectively. Additionally, a Visual Encoder is utilized to extract visual features. Afterward, it leverages a Multi-Scale Cross-Attention Module for visual-textual fusion. The resulting visual-textual representations are fed to a Visual Decoder to synthesize talking face frames.
2.2 Text-Driven Talking Face Generation
The exploration of talking face generation based on first-person statements has not been well-established. Initially, models using Variational Autoencoder (VAE) [20, 24, 37, 40] and CRNN [4, 35] are designed to generate video based on text input. However, the generated frames by these models are unrealistic and over-smoothing, especially in long-sequence video generation tasks [40, 26].Text-driven talking face generation has mainly been divided into two directions: using first-person statements or third-person descriptions. Methods driven by third-person descriptions [13, 25, 41] aim to learn the mapping between objective text descriptions and synthesized video. Most relevant approaches using first-person statements are constructed in a two-stage architecture with two separate steps: converting text to speech and then generating faces from the speech [21, 22, 42, 27, 45]. Consequently, one-stage methods that generate talking faces directly from first-person text captions remain unexplored. Addressing this research gap, our model is constructed as a one-stage end-to-end framework that directly leverages reference talking face frames and first-person text caption to generate high-quality talking faces efficiently. Since relevant exploration has not been widely established, our solution aims to bridge first-person text and dynamic face synthesis to provide insightful guidance for future work.
3 Method
We address the challenging problem of generating realistic talking face videos from first-person text descriptions. Formally, given $M$ text caption tokens $X^{t}_{1:M}$ as the complete sentence and the initial $k$ talking face frames $X^{v}_{1:k}$ as the input, our objective is to synthesize the subsequent talking face frames $X^{v}_{k+1:N}$. In this context, $t$ represents the textual content, and $v$ represents the visual content. The text captions $X^{t}_{1:M}$ and the talking face frames $X^{v}_{1:N}$ are aligned to correspond with each other.
To tackle the problem, we present a novel pipeline – FT2TF: First-Person Statement Text-to-Talking Face Generation, as illustrated in Figure 2. We use two specialized text encoders – the Global Emotion Text Encoder captures overall emotional tone from complete text caption $X^{t}_{1:M}$, and the Linguistic Text Encoder encodes semantic compression from the reference text caption $X^{t}_{m+1:M}$ which corresponds to frames $X^{v}_{k+1:N}$. Alongside these, a Visual Encoder is employed to extract visual features from reference frames $X^{v}_{1:k}$. Inspired by [23], we design and utilize a Multi-Scale Cross-Attention Module that integrates Global and Local Cross-Attention modules to align features from two modalities into the same distribution, aiming to ensure a comprehensive visual-textual fusion across varying scales and feature dimensions. The resulting visual-textual representations are fed to a Visual Decoder to synthesize talking face frames $X^{v}_{k+1:N}$. Implementation details and structural insights are provided in the supplementary material.
3.1 Multimodal Encoders
In this section, we explore the multimodal encoders utilized in FT2TF. Our investigation concentrates on the design insights and functionalities of the Visual Encoder, Global Emotion Text Encoder, and Linguistic Text Encoder. As shown in Figure 2, the features output from these three encoders are denoted as $F^{v}$, $F^{t}_{emo}$, and $F^{t}_{ling}$, respectively.
Visual Encoder. The CNN-based Visual Encoder plays a pivotal role in extracting image features and facial details from talking face videos. Considering the properties of ResBlocks to handle the temporal dynamics and variations in visual input without massive data or strong data augmentations [15], the Visual Encoder is built with ResBlocks as the backbone. The extracted visual features, $F^{v}$, will then be integrated with the textual ones through the Multi-Scale Cross-Attention module. Furthermore, Table 5 illustrates the results of various backbones of the Visual Encoder.
Global Emotion Text Encoder. The Global Emotion Text Encoder harnesses the capabilities of the pre-trained Emoberta [19] model. Emoberta serves as the cornerstone for encoding the overarching emotional context embedded within textual descriptions. In light of this, it presents itself as a prime contender for global attention. By utilizing the entire face frames, we can direct our exclusive focus toward text emotions and their face representations.
Linguistic Text Encoder. We adopt to use the pre-trained GPT-Neo [10] as the Linguistic Text Encoder. In our choice to employ this model, we find motivation in the distinctive characteristics of the GPT-Neo, as opposed to the Emoberta [19]. GPT-Neo has undergone training using a vast language dataset, which empowers it with enhanced linguistic compression for every statement conveyed by the talking face. Consequently, we intend to narrow our attention solely to the region of the lips, aligning with the phonetic articulation of each word, as can be seen in Figure 2(c).
(a) Ground Truth talking face frame sequences.
(b) Corresponding attention maps from Global Cross-Attention.
(c) Corresponding attention maps from Local Cross-Attention.
Figure 3: Visualization of Global and Local Cross-Attention. The red areas indicate high attention weights, while the blue areas indicate the opposite.
3.2 Multi-Scale Cross-Attention and Visual Decoder
Given the encoded textual ($F^{t}_{emo}$, $F^{t}_{ling}$) and visual features ($F^{v}$), we design a Multi-Scale Cross-Attention module that integrates Global and Local Cross-Attention modules. Specifically, features extracted from the Visual Encoder, $F^{v}$, are used in conjunction with the textual features from the Global Emotion Text Encoder, $F^{t}_{emo}$, as the input to the Global Cross-Attention. In this context, $F^{v}$ serves as both the key ($K$) and value ($V$), while $F^{t}_{emo}$ functions as the query ($Q$). A similar process occurs with the Local Cross-Attention where $F^{v}$ serves as $K$ and $V$ and the output of Linguistic Text Encoder, $F^{t}_{ling}$, acts as $Q$.
Global Cross-Attention. This module serves as a pivotal component for infusing global emotional cues from the textual description into the video context. It combines the encoded video features $F^{v}$ with the emotional text description ones $F^{t}_{emo}$ and generates the fusion features $F^{v\rightarrow t}_{emo}$. As shown in Figure 2(b), this integration empowers the model to effectively incorporate emotional expression into facial components such as cheeks, eyes, and eyebrows, enabling more expressive and emotionally accurate face synthesis.
Local Cross-Attention. This module enriches the model’s understanding of local language features in the text. Specifically, by fusing the features $F^{v\rightarrow t}_{ling}$ from $F^{v}$ and $F^{t}_{ling}$, it facilitates the generation of localized and context-aware facial expressions and details. Since the linguistic information is mainly presented by mouth movements, the Local Cross-Attention results in the highest weight in the area around the mouth, as shown in Figure 2(c). This highly enhances the model’s ability to capture fine-grained linguistic nuances and improves the overall coherence of the generated face.
Multi-Scale Cross-Attention. The collective use of Global and Local Cross-Attention ensures a balanced fusion of global emotional context and local linguistic features. As corroborated by Section 4, this multimodal fusion mechanism empowers FT2TF to generate talking face videos with natural lip synchronization and emotionally expressive facial details, resulting in more realistic and engaging outputs.
Visual Decoder. The Visual Decoder is constructed with transpose convolution blocks, aligning each convolutional kernel with its counterpart in the Visual Encoder and matching the channel dimensions to those of $F^{v\rightarrow t}_{emo}$ and $F^{v\rightarrow t}_{ling}$. In this way, the output resolution of the Visual Decoder aligns with the model input. The Visual Decoder conducts channel-wise convolution operations on the two text features and visual features, aiming to reconstruct the visual information embedded in the text features.
The output of the Global and Local Cross-Attention ($F^{v\rightarrow t}_{emo}$, $F^{v\rightarrow t}_{ling}$) are concatenated and subsequently input into the Visual Decoder. The Visual Decoder generates faces from the $k+1$ frame to the $N$ frame $O^{tv}_{k+1},...,O^{tv}_{N}$.
3.3 Pipeline Training
Three key loss functions are constructed for the pipeline training.
Generation Loss. The generation loss, $L_{gen}$, measures the pixel-wise discrepancy between the generated frames and the ground truth ones. This loss primarily focuses on supervising the accurate generation of pixel-level details to ensure the fidelity of the synthesized frames, as shown below:
$L_{gen}=\frac{1}{N}\sum_{i=1}^{N}\left\|X^{v}_{k+i}-O^{tv}_{k+i}\right\|_{1}$ (1)
Synthesis Loss. To improve lip synchronization and the naturalness of lip movements [28], the synthesis loss, $L_{syn}$, leverages the audio component by integrating the Mel-spectrum generated by the Face Synthesizer. While it is not explicitly expressed in mathematical terms, $L_{syn}$ involves the alignment of the synthesized talking face frames with the corresponding audio data $A_{k+i}$:
$L_{syn}=\frac{1}{N}\sum_{i=1}^{N}SyncNet(O^{tv}_{k+i},A_{k+i})$ (2)
Discriminator Loss. The discriminator loss [12], $L_{disc}$, aims to guide the encoder-decoder process to produce more realistic and coherent images. It operates by utilizing a discriminator that provides binary predictions $O_{pred}$ based on whether the input is the generated one or the ground truth. Specifically, the corresponding labels $Y_{disc}$ are assigned the label 1 for generated data and 0 for ground truth, yielding a binary cross-entropy loss:
$\begin{split}L_{disc}=-\frac{1}{N}\sum_{i=1}^{N}(&Y_{disc}\cdot\log(O_{pred})\\ &+(1-Y_{disc})\cdot\log(1-O_{pred}))\end{split}$ (3)
Eventually, the total loss $L_{total}$ is expressed as follows:
$L_{total}=\lambda_{1}L_{gen}+\lambda_{2}L_{syn}+\lambda_{3}L_{disc}$ (4)
4 Experiment
Method Dataset Generation Input PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ LPIPSnormal-$\downarrow$ FIDnormal-$\downarrow$ LipLMDnormal-$\downarrow$ CSIMnormal-$\uparrow$
ATVGnet [5] Lip Region A+V 11.55 0.3944 0.5575 223.26 0.27401 0.1020
Wav2Lip [28] Lip Region A+V 27.92 0.8962 0.0741 43.46 0.02003 0.5925
IP_LAP [46] Lip Region A+V+L 32.91 0.9399 0.0303 27.87 0.01293 0.6523
EAMM [17] LRS2 Whole Face A+V+P 15.17 0.4623 0.3398 91.95 0.15191 0.2318
PC-AVS [47] Whole Face A+V+P 15.75 0.4867 0.2802 110.60 0.07569 0.3927
MakeItTalk [49] Whole Face A+V 17.25 0.5562 0.2237 76.57 0.05024 0.5799
FT2TF (Ours) Whole Face T+V 33.20 0.9901 0.0058 22.52 0.02356 0.9642
ATVGnet [5] Lip Region A+V 10.90 0.3791 0.5667 190.49 0.30564 0.1176
Wav2Lip [28] Lip Region A+V 28.45 0.8852 0.0683 49.60 0.02001 0.5909
IP_LAP [46] Lip Region A+V+L 32.97 0.9222 0.0310 29.96 0.01353 0.6385
EAMM [17] LRS3 Whole Face A+V+P 15.37 0.4679 0.3812 108.83 0.17818 0.2689
PC-AVS [47] Whole Face A+V+P 15.60 0.4732 0.3321 115.25 0.10611 0.3537
MakeItTalk [49] Whole Face A+V 17.78 0.5607 0.2788 97.99 0.08432 0.5465
FT2TF (Ours) Whole Face T+V 31.63 0.9895 0.0085 15.97 0.02545 0.9665
Table 1: Comparison with state-of-the-art talking face generation methods on different evaluation metrics on LRS2 and LRS3 Datasets. Bold results indicate the best among all methods, while the results in the gray background indicate the best among whole face generation methods. “V” - Visual Modality. “A” - Audio Modality. “T” - Textual Modality. “L” - Landmark Source. “P” - Pose Source.
4.1 Datasets
We conduct experiments on two widely used and substantial large-scale text-audio-visual datasets: LRS2 [1] and LRS3 [2].
  • LRS2 [1]: The dataset consists of 48,164 video clips collected from BBC outdoor broadcasts, and is split into the train, validation, and test sets containing 45,839, 1,082, and 1,243 video clips, respectively. Each clip is paired with a text caption containing up to 100 characters.
  • LRS3 [2]: The LRS3 dataset is constructed from over 400 hours of 5,594 indoor TED/TEDx shows with 151,819 videos, along with the corresponding text captions. Specifically, it consists of 32,000 train/val utterances and 1,452 test ones.
While the two datasets share a similar name, it is essential to clarify that there is no overlap between them. Videos in these two datasets differ significantly in terms of the scenes captured, collection methods, character actions, environmental conditions, and other factors. Supplementary material provides data pre-processing details.
4.2 Comparison Methods
We conduct a comprehensive evaluation of our method against state-of-the-art approaches for talking face generation. Due to limited research in first-person text-driven talking face generation, we demonstrate the effectiveness and efficiency of our model by comparing it with audio-driven models on LRS2 and LRS3 datasets. To provide a clear and comprehensive assessment, we categorize these models based on two criteria:
  • Generation Task: All methods are categorized into either Lip Region or Whole Face, with the former generating only the lip region and the latter generating the entire face.
  • Modality: All methods are classified based on the input modality, including Audio+Video (A+V), Text+Video (T+V), Audio+Video+Pose/Landmarks (A+V+P/L).
4.3 Evaluation Metrics
We employ multiple evaluation metrics across various criteria, ranging from evaluating video synthesis quality to evaluating facial expression representation. To evaluate the similarity between generated results and ground truth ones, we apply Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM) [38]. In addition, we utilize Learned Perceptual Image Patch Similarity (LPIPS) [44] and Frechet Inception Distance (FID) [14] for assessing feature-level similarity. Considering the identity preservation, we leverage cosine similarity (CSIM) between identity vectors extracted by the ArcFace [8] network. Moreover, aligning with the approach outlined in [39], we measure the normalized lip landmarks distance (LipLMD) between generated and ground truth frames to evaluate lip synchronization quality.
4.4 Comparison with State-of-the-Art
4.4.1 Quantitative Results
As shown in Table 1, FT2TF not only demonstrates the effectiveness compared to existing models but also achieves state-of-the-art performance across multiple evaluation metrics in talking face generation. This is evidenced by the quantitative results obtained on both LRS2 and LRS3 datasets.
Video Quality Enhancement. In terms of video quality evaluation metrics (PSNR, SSIM, and LPIPS), FT2TF reaches the state-of-the-art. Specifically, in comparison with other whole-face generation approaches on LRS2, it demonstrates significant improvements, with PSNR ($33.20$), SSIM ($0.9901$), and LPIPS ($0.0058$) showing an improvement of at least $92.5\%$, $78.0\%$, and $97.4\%$, respectively. FT2TF consistently outperforms lip-specific methods, even those using ground truth for the remaining parts of the face. Similar findings can be observed for the LRS3. These results show that FT2TF-generated videos exhibit the highest quality and are visually the closest to real-world ones.
Data Distribution Alignment. FID represents the alignment quality of data distribution, in which FT2TF achieves the optimal scores of $22.52/15.97$ on LRS2/3 datasets. The significant decrease of FID (e.g. at least $-70.6\%$ on LRS2 compared to the highest performing face-generation models) indicates the best-aligned data distribution between synthesized frames and real-world ones. This clearly demonstrates that our generated videos are the most statistically indistinguishable from the real ones.
Lip Synchronization and Identity Preservation. Although FT2TF is not exclusively designed for lip generation, our lip synchronization results (LipLMD) outperform all face-generation methods and remain competitive with the latest lip-specific approach (IP_LAP [46]). Since our model learns the entire face representation, it preserves facial expressions completely, capturing every detail. Consequently, we reach the best identity preservation as presented by the CSIM score.
Robustness Across Datasets. FT2TF shows robust performance across distinct and non-overlap datasets. Despite the variations between the two datasets outlined in Section 4.1, FT2TF consistently shows its stable and high-quality generation capabilities. This highlights the model’s adaptability and its ability to generalize across diverse video sources.
Efficiency Analysis. Figure 4 compares the total trainable parameters across various models, including both lip-specific and face-generation approaches. Specifically, when compared with the state-of-the-art methods that only learn lip-around features instead of the whole face (IP_LAP [46], Wav2Lip [28]), our model achieves a higher SSIM score while using a competitive number of trainable parameters. Moreover, our model outperforms other cutting-edge face-generation models significantly. Quantitatively, we achieve a $78.01\%$ improvement in SSIM with only $2.47\%$ more trainable parameters than MakeItTalk [49]. In comparison to other face-generation models, we present substantial advantages in terms of both trainable parameters and performance, with an average performance improvement of $122.88\%$ and an average parameter decrease of $68.94\%$.
Figure 4: Efficiency Comparison. Our analysis includes a comparison of SSIM scores against the total trainable parameters for various models, demonstrating our model’s efficiency.
Figure 5: Qualitative comparison with state-of-the-art methods on LRS2 and LRS3. The three models on the left specialize in lip generation, whereas the others are designed to generate entire faces. Our model consistently generates the most detailed and accurate talking faces across diverse roles, genders, and ages under different lighting and head-poses conditions.
Method
Temp
Change
Lip
Sync
Facial
Details
Video
Quality
MakeItTalk [49] 1.83 1.26 1.70 1.87
IP_LAP [46] 3.13 2.90 2.99 2.82
FT2TF (Ours) 4.50 4.58 4.77 4.43
Table 2: User study on the state-of-the-art lip/face generation methods. Mean Option Scores (MOS) are reported based on four aspects: Temp Change: temporal transition smoothness. Lip Sync: lip synchronization accuracy. Facial Details: facial detail naturalness. Video Quality: overall video quality. We conducted this experiment over three different methods.
Figure 6: Qualitative comparison demonstration in continuous talking face frames Generation. Our model demonstrates enhanced capability in generating talking faces with temporally coherent and accurately synchronized lip movements that align with the spoken text.
4.4.2 Qualitative Results
We process all talking face frames using the same face detector [28], enabling a more detailed comparison of facial expressions. Figure 5 shows the comparison of generation quality between our model and relevant methods on LRS2 and LRS3. Our results stand out as the most similar to the ground truth across diverse conditions, including different character roles and identities, various lighting, and face angles. Benefiting from the Global Emotion Text Encoder, Linguistic Text Encoder, and Multi-scale Cross-Attention modules, our model accurately preserves facial expression details, such as teeth, face shape, head pose, skin color, and texture details. Additionally, we also demonstrate the results for multiple video frames, as shown in Figure 6. A temporal comparison with the state-of-the-art face generation method [49] demonstrates our model constructs the most accurate and consistent video frames across the time domain. Besides the visual demonstration, we also report quantitative results of temporal generation in Section 4.5.
4.5 User Study
We conduct a user study on videos synthesized by the state-of-the-art lip/face generation methods. 30 participants with diverse roles, genders, and ages are invited to watch 4 videos from each method and evaluate on a scale of 1 to 5 for four different aspects: temporal transition smoothness, lip synchronization accuracy, facial detail naturalness, and overall video quality. According to the Mean Opinion Scores (MOS) reported in Table 2, our model highly improves the generation quality compared to state-of-the-art lip/face generation methods from all evaluation domains.
Figure 7: Demonstration of text-driven talking face manipulation. Talking face frames of the left and right sides are synthesized in response to two different textual inputs utilizing the same visual source.
4.6 Text-Driven Face Manipulation Analysis
As one of the valuable real-world applications, we conduct experiments on talking face manipulation based on different text inputs. As shown in Figure 7, we observe two different outputs generated from distinct input texts correspondingly. The first input text, “IF YOU WERE WONDERFUL”, results in generating a detailed facial expression characterized by a smiling face, forehead wrinkles, enhanced cheekbones, and more open eyes. These details contribute to a brighter and more cheerful facial expressions that are consistent with the emotional characteristics conveyed by the input text, reflecting happiness and joy. Meanwhile, the second input text – “IF YOU WERE FEAR” generates faces with contrasting expressions, featuring a more serious and somber representation. Specifically, the forehead does not exhibit prominent wrinkles and the cheekbones appear less pronounced. Moreover, the mouth shape aligns with the intense emotion described in the text “FEAR”. The generated faces accurately reflect the emotion of fear, emphasizing the model’s capacity to adapt facial features in response to different emotional contexts. The reality of our model in achieving these remarkable results can be attributed to the integration of the Global Text Emotion Encoder and the Global Cross-Attention mechanism. The Global Text Emotion Encoder captures emotional semantics from the input text, which is then incorporated by the Global Cross-Attention module for the reflection on different regions of the generated faces. This experimental result showcases the model’s potential in real-world applications such as emotion-driven avatar generation, video dubbing, and virtual assistant interfaces.
Method PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
w/o Global CA 23.03 0.9101 40.59 0.8965
w/o Local CA 21.49 0.8975 40.56 0.8925
w/o Both CA 22.73 0.7317 127.38 0.5682
FT2TF (Ours) 33.20 0.9901 22.52 0.9642
Table 3: Ablation study on Cross-Attention mechanisms. We conduct experiments on the Cross-Attention (CA) mechanisms on LRS2.
4.7 Ablation Analysis
We conduct a comprehensive analysis of the ablation studies on the FT2TF to investigate the impact of key components, including the Cross-Attention mechanisms, loss terms, and different visual encoder backbones. The primary goal is to elucidate the model’s performance and the importance of these elements in generating high-quality face images from first-person text statements.
Cross-Attention Mechanisms. The ablation study on the Cross-Attention module is shown in Table 3. By removing either Global or Local Cross-Attention, a significant drop in all the metrics can be observed. Since emotional coherence and linguistic information are enhanced from the Cross-Attention module as discussed in Section 3.2, the model without it fails to generate high-quality and distribution-aligned faces with the person’s identity being preserved.
Method PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
w/o $L_{syn}$ 28.92 0.9389 40.59 0.8965
w/o $L_{disc}$ 19.90 0.8736 25.66 0.9305
FT2TF (Ours) 33.20 0.9901 22.52 0.9642
Table 4: Ablation study on loss terms. Specifically, we conduct ablation study on $L_{syn}$ and $L_{disc}$ on LRS2.
Role of Loss Terms. We delve into the role of specific loss terms, $L_{syn}$ and $L_{disc}$, as shown in Table 4. $L_{syn}$ focuses on improving the naturalness of lip movements and facial expression details, which leads to a significant gain in all the quantitative metrics. $L_{disc}$ guides the model to generate more realistic frames through binary predictions by the discriminator, contributing to accurate and expressive faces. It is also reflected obviously in the table.
Backbone PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
ViT-B 31.30 0.9690 27.61 0.9378
ViT-L 32.89 0.9690 20.60 0.9434
ResBlocks 33.20 0.9901 22.52 0.9642
Table 5: Ablation study on Visual Encoder backbones. We construct the model with ViT/ResBlocks as the encoder backbone on LRS2.
Encoder Backbones. We investigate the effectiveness of different backbone structures of the Visual Encoder, including ViT with different scales (ViT-B/L) and ResBlocks. As shown in Table 5, our model achieves stable and effective results with various encoder backbones, which proves the flexibility of our pipeline structure. Moreover, since ResBlocks retains information at different scales and inherently captures spatial hierarchies in the visual input, it results in a lead in PSNR, SSIM, and CSIM compared with ViT. Considering the parameter and data efficiency, we utilize ResBlocks as the Visual Encoder backbone.
5 Conclusion
This paper addresses the challenge of generating talking faces through a novel pipeline, FT2TF, which utilizes first-person text captions to replace traditional audio input. Our approach ensures the creation of faces with natural and detailed expressions. By incorporating two large language models, we extract emotional and linguistic features from the text. Moreover, the Multi-Scale Cross-Attention module effectively fuses visual and textual features, contributing to the synthesis of emotionally expressed talking faces. Extensive experiments on the LRS2 and LRS3 datasets underscore the effectiveness of FT2TF, demonstrating its capability to generate realistic, natural, and emotionally expressive talking face from first-person statement text.
[…]
Supplementary Material
A Implementation Details
Video Preprocessing. The facial cropping method is from [28]. To ensure efficient training, we set the input frame size to 96×96.
Audio Preprocessing. Following the methodology presented in [28], we compute Mel-spectrograms from 16kHz audio data using a window size of 800 and a hop size of 200.
Visual Encoder. This module plays a pivotal role in capturing spatial information from the talking face video. It consists of 18 layers of 2D convolutional blocks, each accompanied by Batch Normalization and ReLU activation functions. Additionally, a subset of these modules incorporates residual blocks and skip connections, enhancing its ability to model complex visual features.
Global Emotion Text Encoder. The Global Emotion Text Encoder leverages a pre-trained Emoberta [19] model to encode the overarching emotional tones within the caption text. We import SentenceTransformer from the sentence_transformers library in Python and specifically use the tae898/emoberta-base model.
Linguistic Text Encoder. We adopt a pretrained GPT-Neo [10] model as our Linguistic Text Encoder. The encoder is set up using the transformers library in Python. We import AutoTokenizer and GPTNeoModel from transformers library. The tokenizer is initialized with AutoTokenizer.from_pretrained(”EleutherAI/gpt-neo-2.7B”), and the model is initialized with GPTNeoModel.from_pretrained(”EleutherAI/gpt-neo-2.7B”), respectively.
Visual Decoder. It is composed of six sets of 2D transpose convolution blocks, with each block consisting of a 2D transpose convolution layer and two layers of 2D residual convolution blocks. This architectural design enables the transformation of text descriptions into a coherent and expressive sequence of video frames.
Loss Weight Tuning. $\lambda_{1,2,3}$ are the weight of three losses, fine-tuned for optimal performance. During the initial 300 epochs, we assigned 0.7 to $\lambda_{1}$, 0.09 to $\lambda_{2}$, and 0.21 to $\lambda_{3}$. Subsequently, we adjusted these weights, setting $\lambda_{1}$ to 0.9, $\lambda_{2}$ to 0.03, and $\lambda_{3}$ to 0.07. Additionally, we explore the structural aspects of the Face Synthesizer and the Discriminator, highlighting their respective contributions and characteristics.
Figure S8: Questions in user study. We design 4 questions for each method mentioned in the user study, aiming to evaluate the models from multi-dimensions.
(a) Users’ responses of MakeItTalk [49].
(b) Users’ responses of IP_LAP [46].
(c) Users’ responses of FT2TF (ours).
Figure S9: Example screenshot of users’ responses summary of different methods. The results demonstrate that user evaluations are significantly more positive for FT2TF compared to other methods.
Figure S10: Generation samples using uncurated random samples on LRS2 [1] Test Set. Each pair of rows presents a single sample, with the top row of each pair showing the generated talking face frames and the bottom row showing the ground truth.
Figure S11: Generation samples using uncurated random samples on LRS3 [2] Test Set. Each pair of rows presents a single sample, with the top row of each pair showing the generated talking face frames and the bottom row showing the ground truth.
B Loss Structural Insights
The Face Synthesizer in FT2TF is inspired by the frozen-weight, pretrained Syncnet model [28]. It serves as a critical component in the lip synchronization process by aligning the synthesized talking face frames with the corresponding audio data. While not explicitly detailed through mathematical expressions, the Face Synthesizer integrates audio information to enhance the naturalness of lip movements in the generated talking face frames.
The Discriminator is composed of 14 layers of fully convolutional modules, without the use of Batch Normalization layers or skip connections. It operates as a binary classifier, providing binary predictions ($O_{pred}$) based on whether the input is generated or Ground Truth. These predictions are guided by the binary labels ($Y_{disc}$). This binary classification loss, as detailed in $L_{disc}$, aids in ensuring the quality and authenticity of the synthesized frames.
In summary, FT2TF’s loss functions, coupled with the structural characteristics of the Face Synthesizer and the Discriminator, facilitate a comprehensive training process. These components work synergistically to improve the pixel-level fidelity, lip synchronization, and overall realism of the generated talking face videos, making FT2TF an effective solution for natural and expressive talking face synthesis.
C Description of Compared Methods
Below is a brief overview of the compared methods:
  • ATVGnet [5] generates the region around the lip (e.g. nose+lip) by constructing the face landmark from input audio signal and reference image.
  • Wav2Lip [28] constructs a lip masking and reconstruction strategy for lip generation, with the input as audio and video frames.
  • IP_LAP [46] is proposed recently and is regarded as the state-of-the-art method on audio-driven talking face generation. It leverages information from source audio, video, and landmarks through a two-stage framework.
  • EAMM [17] uses pose sequence, emotion source, and audio information to implement talking face synthesis and manipulation.
  • PC-AVS [47] focuses on pose-controllable talking head generation by devising a low-dimension pose code.
  • MakeItTalk [49] disentangles input audio into content and speaker domains, and predicts 3D landmarks to animate the facial expressions.
D Ablation Study on Text Encoders
Method PSNRnormal-$\uparrow$ SSIMnormal-$\uparrow$ FIDnormal-$\downarrow$ CSIMnormal-$\uparrow$
swapping TE 24.24 0.9181 46.68 0.8655
FT2TF (Ours) 33.20 0.9901 22.52 0.9642
Table S6: Ablation study on Text Encoders. We conduct an ablation study on swapping two Text Encoders (TE) on LRS2.
In the ablation study outlined in Table S6, we investigate the effects of interchanging Text Encoders introduced in the Method Section. The Global Emotional Text Encoder, initially employing Emoberta [19], is replaced by a Linguistic Text Encoder (GPT-Neo [10]) to more thoroughly capture the global linguistic features in caption text. Conversely, the original Linguistic Text Encoder, utilizing GPT-Neo for linguistic encoding of text corresponding to generated talking face frames, is replaced with the Global Emotional Text Encoder (Emoberta) to solely extract emotional features in the remaining sentence, which corresponds to the talking face frames being generated.
This Text Encoder swapping results in a notable decrease in performance metrics. The GPT-Neo exhibits limited capabilities in conveying emotional expressions or depicting facial details. In contrast, the Emoberta is confined to processing only parts of the sentence corresponding to the generated talking face frames, thus restricting its effectiveness in accurately capturing correct sentence-level emotional nuances. This shortcoming can lead to the extraction of inappropriate emotional cues. As a result, the model faces challenges in synchronizing facial details with the emotional aspects embedded in the ground truth during training, causing sub-optimal performance.
E User Study
Questionnaires.
The user study questionnaire is shown in Figure S8. We design 4 questions for each method to let the participant answer. The questions include 4 different aspects: temporal transition smoothness, lip synchronization accuracy, facial detail naturalness, and overall video quality.
Response From Users.
Part of the responses from users are shown in Figure S9. From the results, we can observe that the users respond with much more positive results of FT2TF compared with other methods.
F Datasets
We conduct experiments on two widely used and substantial large-scale text-audio-visual datasets: LRS2 [1] and LRS3 [2].
  • LRS2 [1]: The dataset consists of 48,164 video clips collected from BBC outdoor broadcasts, and is split into the train, validation, and test sets containing 45,839, 1,082, and 1,243 video clips, respectively. Each clip is paired with a text caption containing up to 100 characters. To ensure uniform sentence lengths during model training, we selectively retained videos from the training and test sets that had a total frame count falling within the range of 30 to 35 frames. Consequently, our training set comprises 6,406 videos, while the test set consists of 255 videos.
  • LRS3 [2]: The LRS3 dataset is constructed from over 400 hours of 5,594 indoor TED/TEDx shows with 151,819 videos, along with the corresponding text captions. Specifically, it consists of 32,000 train/val utterances and 1,452 test ones. Similarly, in line with the LRS2 dataset, we selected videos from the training and test sets within the 30 to 35 frame range. This resulted in a training set comprising 2,262 videos and a test set of 172 videos.
G Side-by-Side Comparison
We present more results of generated talking face video frames using uncurated random selections on the test sets of the LRS2 [1] and LRS3 [2] datasets. These results are displayed alongside the original video sequences for comparison, as shown in Figure S10 and Figure S11. The results showcase that our model can consistently generate detailed and accurate talking faces across diverse roles, genders, and ages under different lighting and head-poses conditions.